Alpha industry required 25,000 units/year of a accesory component which will be used in its main product. The ordering cost is Rs.250 per order and the purchase price per unit is Rs.150. The carrying cost per unit per year is 12% of the purchase price. Find economic order quantity, number of orders per annum and time between two orders.

1
Expert's answer
2021-12-28T13:39:01-0500

"EOQ =" "\sqrt {\frac{2*annual demand*ordering cost}{inventory carrying cost per unit per year}}"


"EOQ = \sqrt{\frac{2*250*25000}{18}}"


EOQ = 833

Number of orders/ annum = "\frac{annual demand}{EOQ}"


Number of orders per annum = "\frac{25000}{833}"

= 30

Time between 2 orders = "\frac{EOQ}{Annual demand}"


Time between 2 orders = "\frac{833}{25000}"


= 0.0333